OpenStudy (anonymous):

Put the equation in quadratic formula form. x^2+3x+7=9

OpenStudy (anonymous):

\[\huge x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (campbell_st):

well set the equation to zero by subtracting 9 from both sides and you get \[x^2 + 3x - 2 = 0\] the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for your question a = 1, b = 3 and c = -2 just substitute and calculate away. hope it helps

OpenStudy (anonymous):

The standard form of a quadratic equation looks like \[ax^2+bx+c = 0\] a, b, and c are known values here as campbell explained, and of coarse x is the unknown.

OpenStudy (shamim):

quadratic equation is as ax^2+bx+c=0

OpenStudy (anonymous):

-3\[\frac{ -3\pm \sqrt{9+6}}{ 2 }\]